OpenStudy (anonymous):

Hey, can someone help me out with this calculus problem? I'm stuck here. lim (sqrt(1+x)-sqrt(1-x))/x as x->0 Do I times the numerator and denominator by sqrt(1+x)-sqrt(1-x)?

OpenStudy (anonymous):

multiply denominator and numerator by sqrt(1+x) + sqrt(1-x)

OpenStudy (anonymous):

Thank you, I'll try that now. So it was just reversing the the negative sign to positive I see.

OpenStudy (anonymous):

The reasoning is like what you learned in elementary algebra: $$(a+b)(a-b)=a^2-b^2$$

OpenStudy (anonymous):

^ Yeah

OpenStudy (anonymous):

We want to get rid of radicals in our numerator to see if it simplifies; so we can multiply $$(\sqrt{1+x}-\sqrt{1-x})(\sqrt{1+x}+\sqrt{1-x})=(\sqrt{1+x})^2-(\sqrt{1-x})^2$$

OpenStudy (anonymous):

Ahh, I see now. Thanks for explaining that